Suppose market demand and supply are given by Qd-300 - 4P and QS 50 3P. The equilibrium price is: Multiple Choice $35 $40 $50 $60.
Answer #1

The correct answer is: c)

Reason: at equilibrium, Qd = Qs

=> 300-4P = -50+3P

=> P* = 350/7 = 50.

Thanks!